OpenStudy (anonymous):

The volume of the solid obtained by rotating the region enclosed by: y=x^2 and y=6x about y=36

OpenStudy (ddcamp):

About which axis?

OpenStudy (anonymous):

rotate about which line?

OpenStudy (anonymous):

x^2 = 6x x = 0,6 use disk method: V = pi ∫(outer radius)² - (inner radius) V = pi ∫(36 - x²)² - (36 - 6x)² dx, from 0 to 6 V = 7776pi/5